OpenStudy (anonymous):

on average, at least 2500 pieces of luggage are lost or misdirected each day by united states airlines. Of these, 98% are located by the airlines within 5 days. From a given day's lost luggage, at least how many pieces of luggage are still lost after 5 days?

OpenStudy (anonymous):

98/100*2500=2450 Pieces of luggage still lost= 2500-2450=50

OpenStudy (anonymous):

i got also 50. .